# AutoHFormer: Efficient Hierarchical Autoregressive Transformer for Time Series Prediction

### v1

@@ II-H Theoretical Guarantees
For each time step $t$, the attention mechanism only considers time steps $t^{\prime}$ within the range $[t-W/2,t+W/2]$. Thus, the number of attention computations per time step is $W$, and the total complexity is:
@@
Theorem II.4.
The generalization error of the proposed model is bounded by $C\cdot\sqrt{\frac{\log(L)}{L}}$, where $C$ is a constant depending on the model’s architecture.
Proof.
Let $\mathcal{L}_{\text{emp}}$ be the empirical loss and $\mathcal{L}_{\text{true}}$ be the true loss. By the Rademacher complexity bound, the generalization error satisfies:
$\mathcal{L}_{\text{true}}\leq\mathcal{L}_{\text{emp}}+C\cdot\sqrt{\frac{\log(L)}{L}},$ (19)
where $C$ depends on the model’s architecture and the complexity of the attention mechanism. This bound ensures that the model generalizes well to unseen data as $L\to\infty$. ∎
@@ III Experiments
TABLE II: The statistics of 8 public datasets.
@@
    How effectively does AutoHFormer capture transient dynamics and temporal patterns compared to baselines?
@@ III-A Experiment Settings
Datasets. To rigorously evaluate the performance of our proposed model, we constructed a comprehensive benchmark comprising nine real-world datasets drawn from established repositories [14, 15]. These datasets encompass multiple application domains, including electricity consumption, the four Electricity Transformer Temperature (ETT) datasets (ETTh1, ETTh2, ETTm1, ETTm2) and traffic flow datasets (PESM04 and PEMS). As widely adopted benchmarks in the research community, these datasets facilitate the investigation of critical challenges in fields ranging from smart grid management to transportation analytics. A detailed statistical summary of each dataset is presented in Table II.
@@
  • NLinear [16] proposes to implement a novel normalization strategy for sequence processing. It first centers the input sequence by subtracting its final value, effectively creating a zero-referenced representation. This normalized input is then projected through a standard linear transformation layer. This design effectively decouples the learning of temporal patterns from absolute value ranges while preserving the original data distribution in the final output.
@@
  • DLinear [16] proposes a simple yet effective approach that decomposes time series into trend and residual components, processing each through dedicated linear layers. Surprisingly, this minimalist architecture surpasses sophisticated transformer variants.
@@
  • Linear [16] adopts a parameter-efficient design that employs weight sharing across all variates within a dataset while deliberately avoiding explicit modeling of spatial correlations.
@@
  • Informer [12] uses a sparse attention mechanism that reduces computation from quadratic to near-linear time, while maintaining accurate sequence modeling.
@@
  • Autoformer [17] introduces an innovative architecture combining series decomposition with an Auto-Correlation mechanism to model temporal dependencies. The design specifically addresses the challenge of long-term dependency modeling while reducing the computational complexity typically associated with standard self-attention mechanisms.
@@
  • The iTransformer architecture [18] introduces inverted attention mechanisms for inter-series dependency modeling, achieving notable improvements in multivariate forecasting tasks. However, its tokenization strategy - flattening the entire temporal sequence through a single MLP layer. These architectural constraints fundamentally limit its capability to capture the hierarchical, non-linear evolution characteristics.
  • TimeMixer [19] introduces a multiscale-mixing perspective for modeling temporal variations. TimeMixer is grounded in a key insight: time series exhibit scale-dependent patterns, where fine-grained scales capture microscopic details, while coarse-grained scales reveal macroscopic trends.
@@
  • PatchTST [4] leverages patching and channel-independent techniques to facilitate the extraction of semantic information from single time steps to multiple time steps within time series data.
@@
TABLE III: We present comprehensive results of AutoHFormer and baselines on the ETTh1, ETTh2, Electricity, Exchange, Weather, and Solar-Energy datasets in the autoregressive setting. The lookback length $L$ is fixed at 336, and the forecast length $T$ varies across 96, 192, 336, and 720. Bold font denotes the best model and underline denotes the second best.
@@
Models AutoHFormer(Ours) PatchTST [4] TimeMixer [19] iTransformer [18] Autoformer [17] Informer [12] Linear [16] DLinear [16] NLinear [16]
Metric MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$
ETTm1 96 0.287 0.344 0.298 0.346 0.302 0.351 0.309 0.361 0.723 0.569 1.293 0.862 0.311 0.354 0.301 0.345 0.308 0.350
192 0.329 0.371 0.339 0.374 0.347 0.377 0.347 0.385 0.692 0.549 1.328 0.919 0.345 0.374 0.336 0.376 0.345 0.372
336 0.363 0.392 0.369 0.395 0.401 0.414 0.386 0.407 0.727 0.523 1.483 0.963 0.377 0.394 0.371 0.397 0.380 0.393
720 0.422 0.426 0.427 0.430 0.459 0.445 0.448 0.444 0.773 0.579 1.667 1.014 0.432 0.427 0.426 0.429 0.434 0.428
Avg 0.350 0.383 0.358 0.386 0.377 0.396 0.372 0.399 0.728 0.555 1.442 0.939 0.366 0.387 0.360 0.386 0.366 0.385
Improved - - 2.28% 0.78% 7.71% 3.39% 6.26% 4.17% 108% 44.90% 226.28% 145.16% 4.57% 1.04% 2.85% 0.78% 4.57% 0.52%
ETTm2 96 0.172 0.259 0.170 0.260 0.172 0.262 0.182 0.276 0.277 0.349 0.726 0.648 0.190 0.278 0.171 0.267 0.173 0.261
192 0.236 0.308 0.238 0.305 0.249 0.309 0.243 0.315 0.306 0.365 0.632 0.627 0.249 0.329 0.239 0.320 0.241 0.324
336 0.289 0.342 0.294 0.343 0.360 0.365 0.290 0.344 0.343 0.385 0.683 0.680 0.323 0.381 0.312 0.372 0.291 0.344
720 0.379 0.398 0.380 0.399 0.437 0.413 0.384 0.399 0.433 0.437 0.645 0.589 0.450 0.458 0.445 0.455 0.383 0.417
Avg 0.269 0.326 0.270 0.327 0.304 0.336 0.274 0.333 0.339 0.384 0.671 0.645 0.303 0.361 0.291 0.354 0.271 0.336
Improved - - 0.37% 0.30% 13.01% 3.06% 1.85% 2.14% 26.02% 17.79% 149.44% 97.85% 12.63% 10.74% 8.17% 8.59% 0.74% 3.07%
ETTh1 96 0.382 0.407 0.374 0.400 0.384 0.410 0.398 0.418 0.608 0.529 1.178 0.826 0.431 0.441 0.389 0.412 0.426 0.438
192 0.427 0.436 0.417 0.420 0.432 0.441 0.448 0.453 0.519 0.492 1.163 0.830 0.462 0.459 0.428 0.434 0.458 0.454
336 0.431 0.442 0.457 0.451 0.450 0.453 0.465 0.468 0.643 0.541 1.161 0.819 0.472 0.471 0.447 0.452 0.470 0.464
720 0.464 0.480 0.587 0.529 0.608 0.546 0.547 0.533 1.044 0.710 1.428 0.928 0.506 0.516 0.484 0.499 0.473 0.481
Avg 0.426 0.441 0.458 0.450 0.468 0.462 0.464 0.468 0.703 0.568 1.232 0.850 0.467 0.471 0.436 0.449 0.459 0.459
Improved - - 7.51% 2.04% 9.85% 4.76% 8.92% 6.12% 65.02% 28.79% 189.20% 92.74% 9.62% 6.80% 2.34% 5.39% 7.74% 4.08%
ETTh2 96 0.287 0.350 0.294 0.353 0.297 0.358 0.309 0.363 0.371 0.417 0.901 0.759 0.344 0.397 0.324 0.381 0.294 0.354
192 0.353 0.390 0.374 0.404 0.354 0.394 0.389 0.412 0.426 0.453 0.631 0.629 0.435 0.454 0.416 0.440 0.357 0.395
336 0.347 0.396 0.357 0.399 0.348 0.404 0.372 0.410 0.363 0.465 0.579 0.608 0.491 0.492 0.471 0.480 0.349 0.401
720 0.399 0.436 0.384 0.425 0.411 0.441 0.446 0.462 0.491 0.506 0.734 0.662 0.781 0.629 0.747 0.614 0.403 0.441
Avg 0.346 0.393 0.353 0.395 0.352 0.399 0.379 0.411 0.412 0.460 0.711 0.664 0.512 0.493 0.489 0.478 0.350 0.397
Improved - - 2.02% 0.50% 1.73% 1.52% 9.53% 4.58% 19.07% 17.04% 105.49% 68.95% 47.97% 25.44% 41.32% 21.62% 1.15% 1.02%
Electricity 96 0.127 0.222 0.132 0.226 0.161 0.272 0.151 0.253 0.251 0.364 1.652 1.021 0.165 0.271 0.159 0.264 0.175 0.280
192 0.138 0.246 0.148 0.261 0.183 0.293 0.168 0.267 0.267 0.373 1.518 0.975 0.174 0.281 0.168 0.274 0.185 0.290
336 0.160 0.253 0.165 0.258 0.200 0.312 0.191 0.292 1.054 0.783 1.013 0.827 0.191 0.297 0.186 0.219 0.203 0.305
720 0.200 0.287 0.202 0.290 0.226 0.330 0.232 0.326 1.737 1.068 0.986 0.818 0.225 0.327 0.220 0.322 0.242 0.334
Avg 0.156 0.252 0.161 0.258 0.192 0.301 0.185 0.284 0.827 0.647 1.292 0.910 0.188 0.294 0.183 0.269 0.201 0.302
Improved - - 3.21% 2.38% 23.07% 19.44% 18.58% 12.69% 430.12% 156.74% 728.20% 261.11% 20.51% 16.67% 17.30% 6.74% 28.84% 19.84%
Weather 96 0.153 0.201 0.150 0.198 0.181 0.230 0.159 0.210 0.261 0.325 0.830 0.665 0.177 0.237 0.175 0.234 0.182 0.233
192 0.199 0.246 0.194 0.241 0.217 0.264 0.204 0.250 0.289 0.340 0.687 0.612 0.218 0.275 0.216 0.273 0.225 0.268
336 0.247 0.282 0.248 0.282 0.272 0.302 0.255 0.288 0.340 0.391 0.843 0.677 0.263 0.312 0.261 0.310 0.272 0.302
720 0.317 0.332 0.318 0.334 0.359 0.360 0.321 0.335 0.377 0.401 0.934 0.893 0.322 0.362 0.321 0.361 0.338 0.349
Avg 0.229 0.265 0.227 0.263 0.257 0.289 0.234 0.270 0.316 0.339 0.825 0.711 0.245 0.296 0.243 0.296 0.254 0.288
Improved - - -0.87% -0.75% 12.22% 9.05% 2.18% 1.88% 37.99% 27.92% 260.26% 168.30% 6.99% 11.70% 6.11% 0.699% 10.92% 8.68%
PEMS04 12 0.072 0.173 0.080 0.185 0.073 0.178 0.081 0.186 0.594 0.613 0.252 0.376 0.125 0.244 0.096 0.204 0.104 0.209
24 0.083 0.187 0.103 0.204 0.086 0.197 0.098 0.207 0.527 0.580 0.670 0.591 0.152 0.268 0.126 0.236 0.136 0.250
48 0.101 0.205 0.131 0.229 0.125 0.239 0.126 0.234 0.845 0.763 1.353 0.871 0.191 0.299 0.169 0.273 0.184 0.287
96 0.124 0.225 0.150 0.262 0.196 0.302 0.154 0.263 1.149 0.875 0.949 0.738 0.224 0.322 0.204 0.300 0.227 0.316
Avg 0.095 0.198 0.116 0.220 0.120 0.229 0.114 0.222 0.778 0.707 0.806 0.644 0.173 0.283 0.148 0.253 0.171 0.265
Improved - - 22.11% 11.11% 26.31% 15.65% 20.00% 12.12% 718.95% 257.07% 748.42% 225.25% 82.11% 42.93% 55.79% 27.78% 80.00% 33.84%
PEMS08 12 0.066 0.161 0.074 0.177 0.080 0.175 0.081 0.178 0.769 0.686 0.405 0.426 0.139 0.246 0.103 0.207 0.105 0.213
24 0.082 0.175 0.097 0.199 0.104 0.193 0.108 0.197 0.885 0.751 0.786 0.629 0.183 0.275 0.151 0.243 0.154 0.250
48 0.111 0.192 0.140 0.227 0.149 0.212 0.149 0.225 0.959 0.759 1.293 0.842 0.271 0.319 0.242 0.294 0.254 0.301
96 0.157 0.211 0.201 0.254 0.229 0.246 0.209 0.253 1.226 0.910 1.023 0.750 0.352 0.353 0.328 0.333 0.367 0.341
Avg 0.104 0.185 0.128 0.214 0.120 0.206 0.136 0.213 0.959 0.776 0.646 0.661 0.223 0.298 0.206 0.269 0.220 0.276
Improved - - 23.08% 15.68% 15.38% 11.35% 30.77% 15.14% 822.12% 319.46% 521.15% 257.30% 114.42% 61.08% 98.08% 45.41% 111.54% 49.19%
$1^{st}/2^{nd}$ 34/5 34/5 6/22 6/19 0/4 0/8 0/2 0/0 0/0 0/0 0/0 0/0 0/0 0/1 0/5 0/2 0/2 0/5
@@
Parameter Settings. To be fair comparison, all methods are implemented in an autoregressive mechanism following their GitHub settings. And all codes are shown in the following link https://github.com/lizzyhku/Autotime. For all methods, the learning rate is set as $41e-$ 4. The number of heads is 3. And dimension of keys and values of Transformer is set as 128. The hidden dimension is set as 64. The default setting of window size of our method is 32 at which our method achieves the best performance. The patch length is set as 16. The length of label in the training is set as 48. The sequence length is set as 336 as default.

### v2

@@ II-H Theoretical Guarantees
For each time step $t$, the attention mechanism only considers time steps $t^{\prime}$ within the range [tW/2,t+W/2][t-W/2,t+W/2]. Thus, the number of attention computations per time step is $W$, and the total complexity is:
@@
Theorem II.4 (Generalization Error Bound, Extended).
The generalization error $\mathcal{L}_{\text{true}}$ of AutoHFormer is bounded by:
$\mathcal{L}_{\text{true}}\leq\mathcal{L}_{\text{emp}}+\underbrace{\left(\frac{2}{\sqrt{L}}+\sqrt{\frac{\log(L)}{L}}\right)}_{\text{Rademacher Term}}\cdot C(d,W),$ (16)
where $C(d,W)$ depends on embedding dimension $d$ and window size $W$.
Extended Proof.
We formalize the bound via Rademacher complexity $\mathfrak{R}$: 1) Model Class: AutoHFormer’s hypothesis class $\mathcal{H}$ consists of $d$-dimensional embeddings with $W$-sized causal windows. 2) Rademacher Term: For sequence length $L$, the complexity term follows from:
$\mathfrak{R}(\mathcal{H})=\mathbb{E}\left[\sup_{h\in\mathcal{H}}\frac{1}{L}\sum_{i=1}^{L}\sigma_{i}h(z_{i})\right]\leq\frac{2}{\sqrt{L}}+\sqrt{\frac{\log(L)}{L}},$ (17)
where $\sigma_{i}$ are Rademacher random variables. 3) Architecture Impact: The constant $C(d,W)$ comprises the attention head dimension $d$ (with $|Q/K|_{2}\leq\sqrt{d}$) and the window-restricted complexity $\mathcal{O}(LW)$ from dynamic masking. 4) Generalization: Combining McDiarmid’s inequality with the Rademacher bound yields the theorem’s claim [yin2019rademacher]. ∎
TABLE II: Performance comparison in terms of Pearson correlation
Models Metric ETTm1 ETTm2 ETTh1 ETTh2
AutoHFormer 96 0.844 0.940 0.813 0.900
192 0.821 0.919 0.782 0.871
336 0.803 0.896 0.757 0.869
720 0.768 0.865 0.753 0.859
Avg 0.809 0.905 0.776 0.874
PatchTST [huang2024long] 96 0.842 0.939 0.805 0.898
192 0.817 0.917 0.774 0.870
336 0.802 0.898 0.748 0.868
720 0.770 0.864 0.739 0.857
Avg 0.807 0.904 0.766 0.873
S-Mamba [wang2025mamba] 96 0.834 0.938 0.808 0.899
192 0.809 0.915 0.776 0.869
336 0.792 0.894 0.749 0.868
720 0.754 0.863 0.745 0.858
Avg 0.797 0.902 0.769 0.873
@@ III Experiments
TABLE III: The statistics of 8 public datasets.
@@
    How effectively does AutoHFormer capture transient dynamics and temporal patterns including some edge cases and potential limitations?
@@ III-A Experiment Settings
Datasets. To rigorously evaluate the performance of our proposed model, we constructed a comprehensive benchmark comprising nine real-world datasets drawn from established repositories [haoyietal-informerEx-2023, haoyietal-informer-2021]. These datasets encompass multiple application domains, including electricity consumption, the four Electricity Transformer Temperature (ETT) datasets (ETTh1, ETTh2, ETTm1, ETTm2) and traffic flow datasets (PESM04 and PEMS). As widely adopted benchmarks in the research community, these datasets facilitate the investigation of critical challenges in fields ranging from smart grid management to transportation analytics. A detailed statistical summary of each dataset is presented in Table III.
@@
  • NLinear [zeng2023transformers] proposes to implement a novel normalization strategy for sequence processing.
@@
  • DLinear [zeng2023transformers] proposes a simple yet effective approach that decomposes time series into trend and residual components.
@@
  • Linear [zeng2023transformers] adopts a parameter-efficient design that employs weight sharing across all variates within a dataset while deliberately avoiding explicit modeling of spatial correlations.
@@
  • Informer [zhou2021informer] uses a sparse attention mechanism that reduces computation from quadratic to near-linear time, while maintaining accurate sequence modeling.
@@
  • Autoformer [wu2021autoformer] introduces a architecture combining series decomposition with an Auto-Correlation mechanism to model temporal dependencies.
@@
  • The iTransformer architecture [liu2023itransformer] introduces inverted attention mechanisms for inter-series dependency modeling, achieving notable improvements in multivariate forecasting tasks.
  • TimeMixer [wang2024timemixer] introduces a multiscale-mixing perspective for modeling temporal variations.
@@
  • PatchTST [huang2024long] leverages patching and channel-independent techniques to facilitate the extraction of semantic information from single time steps to multiple time steps within time series data.
  • S-Mamba [wang2025mamba] tokenizes time points for each variate using a linear layer, employs a bidirectional Mamba layer to capture inter-variate correlations, and incorporates a Feed-Forward Network to learn temporal dependencies.
  • Samformer [ilbert2024samformer] proposes a shallow and lightweight transformer model that avoids bad local minima.
  • TimeFilter [DBLP:journals/corr/abs-2501-13041] enables adaptive, fine-grained dependency modeling for time series analysis. The system transforms input sequences into graph representations to capture spatial-temporal relationships.
@@
TABLE IV: We present comprehensive results of AutoHFormer and baselines on the ETTh1, ETTh2, Electricity, Exchange, Weather, and Solar-Energy datasets in the autoregressive setting. The lookback length $L$ is fixed at 336, and the forecast length $T$ varies across 96, 192, 336, and 720. Bold font denotes the best model and underline denotes the second best.
@@
Models AutoHFormer(Ours) PatchTST [huang2024long] S-Mamba [wang2025mamba] Samformer [ilbert2024samformer] TimeFilter [DBLP:journals/corr/abs-2501-13041] TimeMixer [wang2024timemixer] iTransformer [liu2023itransformer] Autoformer [wu2021autoformer] Informer [zhou2021informer] Linear [zeng2023transformers] DLinear [zeng2023transformers] NLinear [zeng2023transformers]
Metric MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$ MSE $\downarrow$ MAE $\downarrow$
ETTm1 96 0.287 0.344 0.298 0.346 0.304 0.359 0.308 0.350 0.290 0.347 0.302 0.351 0.309 0.361 0.723 0.569 1.293 0.862 0.311 0.354 0.301 0.345 0.308 0.350
192 0.329 0.371 0.339 0.374 0.353 0.391 0.343 0.377 0.337 0.375 0.347 0.377 0.347 0.385 0.692 0.549 1.328 0.919 0.345 0.374 0.336 0.376 0.345 0.372
336 0.363 0.392 0.369 0.395 0.386 0.409 0.377 0.395 0.377 0.397 0.401 0.414 0.386 0.407 0.727 0.523 1.483 0.963 0.377 0.394 0.371 0.397 0.380 0.393
720 0.422 0.426 0.427 0.430 0.442 0.442 0.432 0.429 0.428 0.429 0.459 0.445 0.448 0.444 0.773 0.579 1.667 1.014 0.432 0.427 0.426 0.429 0.434 0.428
Avg 0.350 0.383 0.358 0.386 0.371 0.400 0.369 0.396 0.359 0.387 0.377 0.396 0.372 0.399 0.728 0.555 1.442 0.939 0.366 0.387 0.360 0.386 0.366 0.385
Improved - - 2.28% 0.78% 6.00% 4.43% 5.42% 3.39% 2.57% 1.04% 7.71% 3.39% 6.26% 4.17% 108% 44.90% 226.28% 145.16% 4.57% 1.04% 2.85% 0.78% 4.57% 0.52%
ETTm2 96 0.172 0.259 0.170 0.260 0.181 0.274 0.174 0.261 0.173 0.261 0.172 0.262 0.182 0.276 0.277 0.349 0.726 0.648 0.190 0.278 0.171 0.267 0.173 0.261
192 0.236 0.308 0.238 0.305 0.239 0.309 0.243 0.313 0.240 0.309 0.249 0.309 0.243 0.315 0.306 0.365 0.632 0.627 0.249 0.329 0.239 0.320 0.241 0.324
336 0.289 0.342 0.294 0.343 0.296 0.345 0.297 0.349 0.295 0.344 0.360 0.365 0.290 0.344 0.343 0.385 0.683 0.680 0.323 0.381 0.312 0.372 0.291 0.344
720 0.379 0.398 0.380 0.399 0.381 0.401 0.383 0.402 0.393 0.408 0.437 0.413 0.384 0.399 0.433 0.437 0.645 0.589 0.450 0.458 0.445 0.455 0.383 0.417
Avg 0.269 0.326 0.270 0.327 0.274 0.332 0.274 0.331 0.275 0.330 0.304 0.336 0.274 0.333 0.339 0.384 0.671 0.645 0.303 0.361 0.291 0.354 0.271 0.336
Improved - - 0.37% 0.30% 1.85% 1.84% 1.85% 1.53% 2.23% 1.22% 13.01% 3.06% 1.85% 2.14% 26.02% 17.79% 149.44% 97.85% 12.63% 10.74% 8.17% 8.59% 0.74% 3.07%
ETTh1 96 0.382 0.407 0.374 0.400 0.400 0.422 0.387 0.411 0.384 0.409 0.384 0.410 0.398 0.418 0.608 0.529 1.178 0.826 0.431 0.441 0.389 0.412 0.426 0.438
192 0.427 0.436 0.417 0.420 0.436 0.443 0.438 0.440 0.429 0.437 0.432 0.441 0.448 0.453 0.519 0.492 1.163 0.830 0.462 0.459 0.428 0.434 0.458 0.454
336 0.431 0.442 0.457 0.451 0.443 0.452 0.449 0.451 0.440 0.453 0.450 0.453 0.465 0.468 0.643 0.541 1.161 0.819 0.472 0.471 0.447 0.452 0.470 0.464
720 0.464 0.480 0.587 0.529 0.503 0.506 0.476 0.484 0.475 0.483 0.608 0.546 0.547 0.533 1.044 0.710 1.428 0.928 0.506 0.516 0.484 0.499 0.473 0.481
Avg 0.426 0.441 0.458 0.450 0.445 0.455 0.437 0.446 0.432 0.445 0.468 0.462 0.464 0.468 0.703 0.568 1.232 0.850 0.467 0.471 0.436 0.449 0.459 0.459
Improved - - 7.51% 2.04% 4.46% 3.17% 2.34% 1.13% 1.40% 0.91% 9.85% 4.76% 8.92% 6.12% 65.02% 28.79% 189.20% 92.74% 9.62% 6.80% 2.34% 5.39% 7.74% 4.08%
ETTh2 96 0.287 0.350 0.294 0.353 0.300 0.357 0.295 0.354 0.295 0.356 0.297 0.358 0.309 0.363 0.371 0.417 0.901 0.759 0.344 0.397 0.324 0.381 0.294 0.354
192 0.353 0.390 0.374 0.404 0.362 0.406 0.355 0.395 0.357 0.396 0.354 0.394 0.389 0.412 0.426 0.453 0.631 0.629 0.435 0.454 0.416 0.440 0.357 0.395
336 0.347 0.396 0.357 0.399 0.384 0.405 0.349 0.403 0.354 0.405 0.348 0.404 0.372 0.410 0.363 0.465 0.579 0.608 0.491 0.492 0.471 0.480 0.349 0.401
720 0.399 0.436 0.384 0.425 0.409 0.440 0.401 0.439 0.403 0.439 0.411 0.441 0.446 0.462 0.491 0.506 0.734 0.662 0.781 0.629 0.747 0.614 0.403 0.441
Avg 0.346 0.393 0.353 0.395 0.363 0.402 0.350 0.397 0.352 0.399 0.352 0.399 0.379 0.411 0.412 0.460 0.711 0.664 0.512 0.493 0.489 0.478 0.351 0.397
Improved - - 2.02% 0.50% 4.91% 2.29% 1.15% 1.01% 1.73% 1.52% 1.73% 1.52% 9.53% 4.58% 19.07% 17.04% 105.49% 68.95% 47.97% 25.44% 41.32% 21.62% 1.15% 1.02%
Electricity 96 0.127 0.222 0.132 0.226 0.147 0.249 0.164 0.267 0.202 0.283 0.161 0.272 0.151 0.253 0.251 0.364 1.652 1.021 0.165 0.271 0.159 0.264 0.175 0.280
192 0.138 0.246 0.148 0.261 0.167 0.270 0.172 0.276 0.210 0.286 0.183 0.293 0.168 0.267 0.267 0.373 1.518 0.975 0.174 0.281 0.168 0.274 0.185 0.290
336 0.160 0.253 0.165 0.258 0.189 0.291 0.191 0.292 0.223 0.295 0.200 0.312 0.191 0.292 1.054 0.783 1.013 0.827 0.191 0.297 0.186 0.219 0.203 0.305
720 0.200 0.287 0.202 0.290 0.232 0.327 0.229 0.322 0.254 0.328 0.226 0.330 0.232 0.326 1.737 1.068 0.986 0.818 0.225 0.327 0.220 0.322 0.242 0.334
Avg 0.156 0.252 0.161 0.258 0.183 0.284 0.189 0.289 0.222 0.298 0.192 0.301 0.185 0.284 0.827 0.647 1.292 0.910 0.188 0.294 0.183 0.269 0.201 0.302
Improved - - 3.21% 2.38% 17.30% 12.69% 21.15% 14.68% 42.3% 18.2% 23.07% 19.44% 18.58% 12.69% 430.12% 156.74% 728.20% 261.11% 20.51% 16.67% 17.30% 6.74% 28.84% 19.84%
Weather 96 0.153 0.201 0.150 0.198 0.157 0.207 0.174 0.244 0.183 0.226 0.181 0.230 0.159 0.210 0.261 0.325 0.830 0.665 0.177 0.237 0.175 0.234 0.182 0.233
192 0.199 0.246 0.194 0.241 0.204 0.250 0.217 0.259 0.229 0.264 0.217 0.264 0.204 0.250 0.289 0.340 0.687 0.612 0.218 0.275 0.216 0.273 0.225 0.268
336 0.247 0.282 0.248 0.282 0.253 0.288 0.264 0.294 0.279 0.300 0.272 0.302 0.255 0.288 0.340 0.391 0.843 0.677 0.263 0.312 0.261 0.310 0.272 0.302
720 0.317 0.332 0.318 0.334 0.325 0.339 0.329 0.339 0.348 0.346 0.359 0.360 0.321 0.335 0.377 0.401 0.934 0.893 0.322 0.362 0.321 0.361 0.338 0.349
Avg 0.229 0.265 0.227 0.263 0.234 0.271 0.246 0.298 0.259 0.284 0.257 0.289 0.234 0.270 0.316 0.339 0.825 0.711 0.245 0.296 0.243 0.296 0.254 0.288
Improved - - -0.87% -0.75% 2.18% 2.26% 7.43% 12.45% 13.10% 7.16% 12.22% 9.05% 2.18% 1.88% 37.99% 27.92% 260.26% 168.30% 6.99% 11.70% 6.11% 0.699% 10.92% 8.68%
PEMS04 12 0.072 0.173 0.080 0.185 0.083 0.190 0.142 0.260 0.115 0.228 0.073 0.178 0.081 0.186 0.594 0.613 0.252 0.376 0.125 0.244 0.096 0.204 0.104 0.209
24 0.083 0.187 0.103 0.204 0.100 0.210 0.173 0.284 0.133 0.241 0.086 0.197 0.098 0.207 0.527 0.580 0.670 0.591 0.152 0.268 0.126 0.236 0.136 0.250
48 0.101 0.205 0.131 0.229 0.124 0.234 0.214 0.314 0.157 0.264 0.125 0.239 0.126 0.234 0.845 0.763 1.353 0.871 0.191 0.299 0.169 0.273 0.184 0.287
96 0.124 0.225 0.150 0.262 0.145 0.255 0.247 0.335 0.171 0.279 0.196 0.302 0.154 0.263 1.149 0.875 0.949 0.738 0.224 0.322 0.204 0.300 0.227 0.316
Avg 0.095 0.198 0.116 0.220 0.113 0.222 0.194 0.298 0.144 0.253 0.120 0.229 0.114 0.222 0.778 0.707 0.806 0.644 0.173 0.283 0.148 0.253 0.171 0.265
Improved - - 22.11% 11.11% 18.94% 12.12% 104.21% 50.50% 51.57% 27.78% 26.31% 15.65% 20.00% 12.12% 718.95% 257.07% 748.42% 225.25% 82.11% 42.93% 55.79% 27.78% 80.00% 33.84%
PEMS08 12 0.066 0.161 0.074 0.177 0.085 0.183 0.170 0.265 0.091 0.207 0.080 0.175 0.081 0.178 0.769 0.686 0.405 0.426 0.139 0.246 0.103 0.207 0.105 0.213
24 0.082 0.175 0.097 0.199 0.111 0.205 0.218 0.292 0.166 0.273 0.104 0.193 0.108 0.197 0.885 0.751 0.786 0.629 0.183 0.275 0.151 0.243 0.154 0.250
48 0.111 0.192 0.140 0.227 0.155 0.234 0.314 0.333 0.305 0.337 0.399 0.212 0.149 0.225 0.959 0.759 1.293 0.842 0.271 0.319 0.242 0.294 0.254 0.301
96 0.157 0.211 0.201 0.254 0.212 0.260 0.407 0.361 0.353 0.379 0.229 0.246 0.209 0.253 1.226 0.910 1.023 0.750 0.352 0.353 0.328 0.333 0.367 0.341
Avg 0.104 0.185 0.128 0.214 0.140 0.220 0.277 0.312 0.228 0.299 0.120 0.206 0.136 0.213 0.959 0.776 0.646 0.661 0.223 0.298 0.206 0.269 0.220 0.276
Improved - - 23.08% 15.68% 34.62% 18.91% 118.26% 68.64% 119.23% 61.62% 15.38% 11.35% 30.77% 15.14% 822.12% 319.46% 521.15% 257.30% 114.42% 61.08% 98.08% 45.41% 111.54% 49.19%
$1^{st}/2^{nd}$ 34/5 34/5 6/21 6/19 0/1 0/0 1/0 0/0 0/3 0/1 0/4 0/8 0/1 0/0 0/0 0/0 0/0 0/0 0/0 0/1 0/3 0/1 0/1 0/5
@@
Parameter Settings. To be fair comparison, all methods are implemented in an autoregressive mechanism following their GitHub settings. And all codes are shown in the following link https://github.com/CoderPowerBeyond/AutoHFormer. For all methods, the learning rate is set as 1e41e-4. The number of heads is 3. And dimension of keys and values of Transformer is set as 128. The hidden dimension is set as 64. The default setting of window size of our method is 32 at which our method achieves the best performance. The patch length is set as 16. The length of label in the training is set as 48. The sequence length is set as 336 as default.